If the Ka of a monoprotic weak acid is 6.8 × 10−6, what is the pH of a 0.11 M solution of this acid?
Answer #1

pH of the solution of a weak acid = 1/2 pKa - 1/2 log C where pKa = - log Ka and C = concentration.

Given, Ka = 6.8 x 10-6 and C = 0.11 M

So, pKa = 5.167

and pH = (5.167/2) - (log(0.11)/2)

   = 2.58 - (-0.47) = 3.05